# Attention Normalization Impacts Cardinality Generalization in Slot Attention

### v1

@@ 6 Related Work
Previous works have discussed generalization capabilities of object-centric representations (Dittadi et al., 2022; Seitzer et al., 2023). That the number of slot latents has influence on model performance has been noted in (Seitzer et al., 2023) where the authors illustrate that varying the number of slots has a significant impact on segmentation quality, determining whether objects are split into constituent parts or discovered in their entirety. To date, few works investigate the role of attention normalization in the performance of Slot Attention. The first ablation evaluated for Slot Attention in the original paper (Locatello et al., 2020) is almost identical to the approach we discuss in Subsection 4.1, differing from our method crucially in that the weighted sums are not scaled by a constant, which appears to result in poor training behavior. In a second ablation, the authors modify the first ablation by normalizing the update codes via layer normalization (Ba et al., 2016). Comparable performance is reported to the original method. Normalization in Slot Attention has been investigated in (Zhang et al., 2023), where the authors use the Sinkhorn-Knopp iteration to normalize attention matrices. In contrast to our work, they do not investigate the impact on generalization capabilities and substantially increase the complexity of the SA module.
@@ Appendix A Reconstruction Loss
We first consider the experiment in which we trained a model with 7 slots on CLEVR6. We note in Subfigures 6(a) and 6(b) that all normalization approaches suffer under excess slots during inference. The layer norm variant generally performs best in this context, while the baseline and weighted sum variants perform comparably to each other. In Figure 6(c), we note that reconstruction quality decreases with increasing object count for all variants. It can be observed that this deterioration progresses fastest for the baseline, slowest with batch normalization, and at a seemingly similar rate for the layer norm and weighted sum variants.
@@
Figure 7 illustrates the behavior of reconstruction losses for models trained on CLEVR6 with excess slots. We find that, analogous to our observations in Subsection 5.1, varying slot count during inference has a lesser effect than in the previous experiment (compare Subfigures 7(a) and 7(b) to Subfigures 6(a) and 6(b)). Generally, the weighted sum variant has the highest reconstruction loss, while the layer norm variant has the lowest. The baseline and the batch norm variant perform comparably. As before, we observe in Subfigure 7(c) that reconstruction loss increases with increasing object count.
@@
We now shift our focus to the Dinosaur models. Figure 8 shows the reconstruction loss for the models trained on MOVi-C10 with 11 slots. Contrary to previous observations, we note in Subfigures 8(a) and 8(c) that reconstruction losses improve as the slot count increases. Generally, it appears that the weighted sum variant performs worst w.r.t. the $\ell^{2}$ loss, while the other variants perform comparably to each other. We observe analogous behavior in Figures 9 and 10.

### v2

@@ 6 Related Work
In this work, we follow a longer tradition of using autoencoders to obtain semantic scene decompositions (Greff et al., 2016; 2017; 2019; Crawford & Pineau, 2019; Burgess et al., 2019; Lin et al., 2020; Engelcke et al., 2020; Locatello et al., 2020). An early work in this area is Neural-EM (Greff et al., 2017), which performs expectation maximization at the image level. It is proceeded by IODINE (Greff et al., 2019) and MONet (Burgess et al., 2019), which learn variational autoencoders. Slot Attention (Locatello et al., 2020) has already been used in many derivative works to scale object-centric learning to increasingly complex datasets. In particular, motion cues (Kipf et al., 2022; Elsayed et al., 2022; Wu et al., 2023), high-level image features (Seitzer et al., 2023) and powerful decoder models (Singh et al., 2022a; b) were found to be useful for obtaining desired behavior on real-world datasets.
Several previous works have discussed generalization capabilities of object-centric representations (Dittadi et al., 2022; Seitzer et al., 2023). That the number of slot latents has influence on model performance has been noted by Seitzer et al. (2023) where the authors illustrate that varying the number of slots has a significant impact on segmentation quality, determining whether objects are split into constituent parts or discovered in their entirety. To date, only few works investigate the role of attention normalization in the performance of Slot Attention. The first ablation evaluated for Slot Attention in the original paper (Locatello et al., 2020) is almost identical to the approach we discuss in Subsection 4.1, differing from our method crucially in that the weighted sums are not scaled by a constant, which appears to result in poor training behavior. In a second ablation, the authors modify the first ablation by normalizing the update codes via layer normalization (Ba et al., 2016). Comparable performance is reported to the original method.
Normalization in Slot Attention has been investigated in (Zhang et al., 2023), where the authors propose to use the Sinkhorn-Knopp iteration to normalize attention matrices. In contrast to our work, they do not investigate the impact on generalization capabilities and substantially increase the complexity of the SA module. More broadly, the use of the Sinkhorn-Knopp algorithm for normalization in multi-head attention modules has been scrutinized before by Sander et al. (2022).
@@ Appendix A Reconstruction Loss
We first consider the experiment in which we trained a model with 7 slots on CLEVR6. We note in Subfigures 7(a) and 7(b) that all normalization approaches suffer under excess slots during inference. The layer norm variant generally performs best in this context, while the baseline and weighted sum variants perform comparably to each other. In Figure 7(c), we note that reconstruction quality decreases with increasing object count for all variants. It can be observed that this deterioration progresses fastest for the baseline, slowest with batch normalization, and at a seemingly similar rate for the layer norm and weighted sum variants.
@@
Figure 8 illustrates the behavior of reconstruction losses for models trained on CLEVR6 with excess slots. We find that, analogous to our observations in Subsection 5.1, varying slot count during inference has a lesser effect than in the previous experiment (compare Subfigures 8(a) and 8(b) to Subfigures 7(a) and 7(b)). Generally, the weighted sum variant has the highest reconstruction loss, while the layer norm variant has the lowest. The baseline and the batch norm variant perform comparably. As before, we observe in Subfigure 8(c) that reconstruction loss increases with increasing object count.
@@
We now shift our focus to the Dinosaur models. Figure 9 shows the reconstruction loss for the models trained on MOVi-C10 with 11 slots. Contrary to previous observations, we note in Subfigures 9(a) and 9(c) that reconstruction losses improve as the slot count increases. Generally, it appears that the weighted sum variant performs worst w.r.t. the $\ell^{2}$ loss, while the other variants perform comparably to each other. We observe analogous behavior in Figures 10 and 11.
@@ Appendix F Technical Details Regarding Object Discovery on MOVi-C
Appendix G Pseudocode
In Algorithms 1 and 2, we illustrate how the weighted sum and batch norm variants differ from the weighted mean variant in pseudo PyTorch code. We illustrate this in a diff format.
Algorithm 1 Diff of Weighted Sum Variant
{minted}
[linenos]diff … bs, N, d_in = inputs.shape k, v = self.key_map(inputs), self.value_map(inputs) for idx in range(num_iters): slots_prev = slots slots = self.norm_slots(slots) q = self.query_map(slots) dots = torch.einsum("bid,bjd->bij", q, k) / np.sqrt(q.size(-1)) attn = dots.softmax(dim=1) - attn = (attn + eps) / (attn + eps).sum(dim=-1, keepdim=True) updates = torch.einsum("bjd,bij->bid", v, attn) + updates = updates / N …
Algorithm 2 Diff of Batch Norm Variant
{minted}
[linenos]diff … bs, N, d_in = inputs.shape k, v = self.key_map(inputs), self.value_map(inputs) + var, mean = None, None for idx in range(num_iters): slots_prev = slots slots = self.norm_slots(slots) q = self.query_map(slots) dots = torch.einsum("bid,bjd->bij", q, k) / np.sqrt(q.size(-1)) attn = dots.softmax(dim=1) - attn = (attn + eps) / (attn + eps).sum(dim=-1, keepdim=True) updates = torch.einsum("bjd,bij->bid", v, attn) + if idx == 0 and self.training: + var, mean = torch.var_mean(updates, correction=0) + self.update_buffers(var, mean) + elif idx == 0 and not self.training: + var, mean = self.var_buffer, self.mean_buffer + updates = self.alpha * (updates - mean) / torch.sqrt(var + eps) + self.beta …
Appendix H Visual Results
In Tables 6 and 7, we present visual results from object discovery on the CLEVR10 and MOVi-C10 datasets, respectively. On CLEVR, we show both reconstructions and segmentations, while we only provide segmentations on MOVi-C, as the reconstructions are high-dimensional ViT features. In both cases, we perform inference with a high slot count (21).
Input Baseline Layer Norm Weighted Sum Batch Norm
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 6: Visual results for object discovery on CLEVR10, trained on CLEVR6 with 7 slots. Showing reconstructions and segmentations. Evaluated with 21 slots.
Baseline Layer Norm Weighted Sum Batch Norm
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 7: Visual results for object discovery on MOViC10, trained on MOViC10 with 11 slots. Evaluated with 21 slots.
Appendix I Property Prediction
We further illustrate the proposed normalizations on a property prediction task on the CLEVR10 dataset. We closely follow the experimental setup detailed by Locatello et al. (2020), using three seeds per variant. We train the models on the CLEVR10 dataset using 10 slots. In Figure 12, we plot how the foreground ARI varies as the number of slots is modified during inference. Consistent with our observations on the object discovery task, we find that the weighted mean and layer norm variants suffer from excess slots, while the proposed normalizations are robust to them. Overall, we find that the batch norm variant performs best w.r.t. foreground segmentation quality.
Figure 12: F-ARI ($\uparrow$) for property prediction on CLEVR10
In Figure 13, we additionally show the mean average prediction at various distance thresholds, as defined in (Locatello et al., 2020). We observe a similar behavior as before, namely that the proposed variants are robust to high slot counts during inference, while the weighted mean and layer norm variants are not. However, we generally find that the weighted mean variant appears to perform best at low slot counts.
(a) mAP@1.0 ($\uparrow$) on CLEVR10
(b) mAP@0.5 ($\uparrow$) on CLEVR10
(c) mAP@0.25 ($\uparrow$) on CLEVR10
(d) mAP@0.125 ($\uparrow$) on CLEVR10
Figure 13: Mean average precision at different distance thresholds for property prediction on CLEVR.
Appendix J Ablation of Normalization Schemes
In this section, we study two alternative normalizations methods. Firstly, we consider the weighted sum variant with the scaling parameter chosen as $C=1$. We refer to this method as "unnormalized". Secondly, we consider a variant which ablates the normalization across the batch axis from the batch normalized variant. I.e., we compute mean and variance for each instance separately across only the slot and layer axes. Hence, we also do not have to keep a moving average of the normalizing statistics for inference. Instead the normalization behaves identically during training and inference. As in the batch normalized variant, two scalar values $\alpha$ and $\beta$ are learned. We refer to this variant as "K-D-Layer Norm", to reflect that we are performing a layer normalization across the slot (K) and layer (D) axes. We perform experiments on the property prediction task on CLEVR with three seeds per variant.
The unnormalized variant fails to obtain object-centric behavior, as becomes apparent from the low foreground ARI across all slot counts, as shown in Figure 14. In Figure 15, we observe that the mean average precision suffers correspondingly. This underscores the importance of scaling the update codes appropriately to achieve competitive performance. While the K-D-Layer Norm performs better, we find that as the other baseline variants, it is not robust to high slot counts. Moreover, on this specific task, it seems to generally underperform in terms of mAP when compared to the other object-centric variants.
Figure 14: F-ARI ($\uparrow$) for property prediction on CLEVR10
(a) mAP@1.0 ($\uparrow$) on CLEVR10
(b) mAP@0.5 ($\uparrow$) on CLEVR10
(c) mAP@0.25 ($\uparrow$) on CLEVR10
(d) mAP@0.125 ($\uparrow$) on CLEVR10
Figure 15: Mean average precision at different distance thresholds for property prediction on CLEVR.